Concord Appliance Center is an experienced home appliance dealer. Concord Appliance Center also offers a number of services for the home appliances that it sells. Assume that Concord Appliance Center sells ovens on a standalone basis. Concord Appliance Center also sells installation services and maintenance services for ovens. However, Concord Appliance Center does not offer installation or maintenance services to customers who buy ovens from other vendors. Pricing for ovens is as follows.

Oven only $800
Oven with installation service 860
Oven with maintenance services 990
Oven with installation and maintenance services 1,010


In each instance in which maintenance services are provided, the maintenance service is separately priced within the arrangement at $190. Additionally, the incremental amount charged by Concord Appliance Center for installation approximates the amount charged by independent third parties. Ovens are sold subject to a general right of return. If a customer purchases an oven with installation and/or maintenance services, in the event Concord Appliance Center does not complete the service satisfactorily, the customer is only entitled to a refund of the portion of the fee that exceeds $800.

Assume that a customer purchases an oven with both installation and maintenance services for $1,010.

(b) Indicate the amount of revenue that should be allocated to the oven, the installation, and to the maintenance contract. (Do not round intermediate calculations. Round final answers to 2 decimal places.)

Oven

$enter a dollar amount rounded to 2 decimal places

Installation

$enter a dollar amount rounded to 2 decimal places

Maintenance

$
Answer #1

Answer:

Step 1)The standalone cost of installation service = cost of oven with installation -cost of oven only

                                            = 860-800

                                             = 60

Step 2)Allocation of purchase cost and weights of standalone price

Item Standalone price Weight Share in purchase cost of 1010
Oven 800 800/1050 = 76.19048% 1010*76.19048%= 769.52
Installation 60 60/1050= 5.71429% 1010*5.71429%= 57.71
Maintenance 190 190/1050 = 18.09524% 1010*18.09524%= 182.76
1050 1010

b)

Item Amount of revenue allocated
Oven 769.52
Installation 57.71
Maintenance 182.76
Total 1010